OpenStudy (anonymous):

solve for x 0=(2^x-4)

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

how is that?

OpenStudy (anonymous):

2^x=4 2 to what power equal 4 the only answer is 2

OpenStudy (anonymous):

x is 2 or -2

OpenStudy (anonymous):

Not -2

OpenStudy (anonymous):

is that possible to make it =4

OpenStudy (anonymous):

That give 2^-2 = 1/4 not 4

OpenStudy (anonymous):

unless you meant to write x^2-4=0, the only solution is +2